The number of valence electrons in transition metals could very in different ceramics. Calculate the number of valence electrons of Nb in Nb oxides NbO, NbO2, and Nb2O5.
Answer #1

Electronic configuration of Nb : [Kr] 4d4 5s1

Number of valence electrons = 5

1. In NbO : oxidation state of Nb = +2

Valence electrons = 5 - 2 = 3

2. In NbO2 : oxidation state of Nb = +4

Valence electrons = 5 - 4 = 1

3. In Nb2O4 : oxidation state of Nb = +4

Valence electrons = 5 - 4 = 1